# Direct Training High-Performance Deep Spiking Neural Networks: A Review of Theories and Methods

### v1

@@ Abstract
Spiking neural networks (SNNs) offer a promising energy-efficient alternative to artificial neural networks (ANNs), in virtue of their high biological plausibility, rich spatial-temporal dynamics, and event-driven computation. The direct training algorithms based on the surrogate gradient method provide sufficient flexibility to design novel SNN architectures and explore the spatial-temporal dynamics of SNNs. According to previous studies, the performance of models is highly dependent on their sizes. Recently, direct training deep SNNs have achieved great progress on both neuromorphic datasets and large-scale static datasets. Notably, transformer-based SNNs show comparable performance with their ANN counterparts. In this paper, we provide a new perspective to summarize the theories and methods for training deep SNNs with high performance in a systematic and comprehensive way, including theory fundamentals, spiking neuron models, advanced SNN models and residual architectures, software frameworks and neuromorphic hardware, applications, and future trends.
@@ 3 Introduction
There are two mainstream pathways to obtain deep SNNs: ANN-to-SNN conversion and direct training through the surrogate gradient method. Firstly, in ANN-to-SNN conversion (Cao et al., 2015; Hunsberger and Eliasmith, 2015; Rueckauer et al., 2017; Bu et al., 2022; Meng et al., 2022; Wang et al., 2022), a pre-trained ANN is converted to an SNN by replacing the ReLU activation layers with spiking neurons and adding scaling operations like weight normalization and threshold balancing. This conversion process suffers from long converting time steps, which causes high computational consumption in practice. In addition, the converted SNNs obtained in this way are constrained by the original ANNs’ architecture, thus the direct exploration of the virtues of SNNs is limited. Secondly, in the field of direct training, SNNs are unfolded over simulation time steps and trained with backpropagation through time (Lee et al., 2016; Shrestha and Orchard, 2018). Due to the non-differentiability of spiking neurons, the surrogate gradient method is employed for backpropagation (Lee et al., 2020b; Neftci et al., 2019; Fang et al., 2021a, b; Zhou et al., 2023c). On one hand, this direct training method can handle temporal data and also achieve decent performance on large-scale static datasets, with only a few time steps. On the other hand, it can provide sufficient flexibility for designing novel architectures specifically for SNNs and exploring the properties of SNNs directly. Therefore, the direct training method has received more attention recently.
@@ 4 Spiking Neuron Models
LIF (Leaky Integrate-and-Fire) neuron is one of the most commonly used neurons in SNNs (Zhou et al., 2023c, a, b), which is simple but retains biological characteristics. The dynamics of LIF are described as:
@@ 5.2.1 Surrogate Gradient
Eq. 13 is the derivative of Eq. 12. In addition, SG could be the derivative of Sigmoid (Zhou et al., 2023c, a, b), tanh (Guo et al., 2022a), or rectangular (Wu et al., 2018, 2019) functions, etc. To address the problem of gradient vanishing caused by surrogate gradient function with fixed parameters, Lian et al. proposed the Learnable Surrogate Gradient (LSG) (Lian et al., 2023), in which a learnable parameter is used to adjust the gradient-available interval.
@@ 5.2.2 Loss Function and Backpropagation
Distinct from ANNs, there’s an additional dimension in SNNs, the temporal domain. For spiking neurons, the membrane potential in the current step depends on the membrane potential in the previous time step, that is, there is a time dependence. Thus, backpropagation in ANNs does not apply to SNNs. Spatio-temporal backpropagation (STBP) (Wu et al., 2018), proposed by Wu et al., takes the gradient update in both the spatial domain and temporal domain into account to train SNNs. Backpropagation Through Time (BPTT) (Bird and Polivoda, 2021), originally developed for recurrent neural networks (RNNs), is applied to SNNs due to their similar characteristics to those of RNNS. BPTT is described as follows:
$\frac{\partial\mathcal{L}}{\partial\mathbf{W}}$ $=\sum_{t}\sum_{t^{\prime}\leq t}\frac{\partial\mathcal{L}[t]}{\partial\mathbf{W}[t^{\prime}]}\frac{\partial\mathbf{W}[t^{\prime}]}{\partial\mathbf{W}}$ (17)
$=\sum_{t}\sum_{t^{\prime}\leq t}\frac{\partial\mathcal{L}[t]}{\partial\mathbf{W}[t^{\prime}]},$ (18)
where $\mathcal{L}[t]$ is the loss function at time $t$, $\mathbf{W}[t^{\prime}]$ and $\mathbf{W}$ are the weight, and $\frac{\partial\mathbf{W}[t^{\prime}]}{\partial\mathbf{W}}=1$ since the weight $\mathbf{W}$ is shared along all timesteps. The combination of BPTT and surrogate gradient in SNNs is common, but it also faces problems such as excessive training memory consumption and the neglect of biological online learning. Therefore, Xiao et al. proposed an online training through time (OTTT) (Xiao et al., 2022) algorithm.
@@ 6.1 Transformer-based Spiking Neural Networks
Biological realism tends to model neural networks with high biological plausibility to simulate the complex biological mechanism of the brain. It often lacks the consideration of computational efficiency and performance optimization on general application tasks. Traditional ANNs often prioritize task performance over biological realism and computational energy consumption. SNNs have great potential to own the characteristics of biological plausibility, low computational energy consumption, and high task performance simultaneously. Especially several direct training Transformer-based SNNs have broken through 80% top-1 accuracy on ImageNet-1K, which instills great optimism in the application of SNNs.
@@ 7.1 Software Frameworks for Training Spiking Neural Networks
One category of frameworks includes brain simulators such as NEURON (Hines and Carnevale, 1997) and Brian (Goodman and Brette, 2009), which not only enhance the scalability and computational efficiency of models but also encompass cognitive functions such as perception, decision-making, and reasoning. The SNNs constructed by these frameworks exhibit a high degree of biological plausibility, making them suitable for studying the functionalities of real neural systems. They support biologically interpretable learning rules such as Spike-Timing-Dependent Plasticity (STDP) (Bi and Poo, 1998), playing a significant role in advancing the field of neuroscience. However, these frameworks lack core computational functionalities required for deep learning, such as automatic differentiation, rendering them incapable of performing machine learning tasks.
@@
Another category of brain-inspired computing frameworks comprises deep spiking computation frameworks. Deep SNNs involve a substantial amount of matrix operations across spatial and temporal dimensions, a variety of neurons, neuromorphic datasets, and deployments on neuromorphic chips. The modeling and application processes are complex, and achieving high-performance acceleration is challenging. To address these issues, spiking deep learning frameworks need to support the construction, training, and deployment of deep SNNs, and be capable of acceleration based on spike operations. Frameworks such as BindsNET (Hazan et al., 2018), NengoDL (Rasmussen, 2019), SpykeTorch (Mozafari et al., 2019), Norse, SpyTorch, SNNTorch (Eshraghian et al., 2023) and SpikingJelly (Fang et al., 2023a) have been developed. They utilize simple spiking neurons to reduce computational complexity, making them suitable for machine learning research. Among them, BindsNet (Hazan et al., 2018) primarily focuses on machine learning and reinforcement learning; NengoDL (Rasmussen, 2019) converts ANNs to obtain deep SNNs but does not support direct training of SNNs using surrogate gradient methods; SpyTorch is a demonstrative framework that only provides basic surrogate gradient examples; SpyTorch (Mozafari et al., 2019) introduces a new type of surrogate gradient method named SuperSpike. These frameworks can implement some simple machine learning and reinforcement learning models, but they still lack deep learning capabilities for SNNs. Norse is attempting to introduce the sparse and event-driven characteristics of SNNs and supports many typical spiking neuron models. It is in the development stage and has not been officially released yet. SNNTorch supports some variants of online backpropagation algorithms that are more biologically plausible and support large-scale SNN computation. SpikingJelly (Fang et al., 2023a) is a full-stack toolkit for preprocessing neuromorphic datasets, building deep SNNs, optimizing their parameters, and deploying SNNs on neuromorphic chips, which shows remarkable extensibility and flexibility, enabling users to accelerate custom models at low costs through multilevel inheritance and semiautomatic code generation. In summary, the development of existing software frameworks is essentially in its early stages, and there is still a long way to go in terms of functionality enhancement and performance optimization.
@@ 7.2 Neuromorphic Hardware for Spiking Neural Networks
7.3 Software and Hardware Interaction
The deployment of algorithms for SNNs onto neuromorphic chips typically requires certain computational and application software frameworks. The computational software frameworks mentioned earlier in the section usually support simulations on mainstream CPUs and GPUs. However, there is no explicit mention of deployment on neuromorphic chips, which remains speculative. And among the previously mentioned neuromorphic chips, a small number are designed in collaboration with application software. In the literature survey conducted on neuromorphic chips, it was found that only 27 percent connected to application softwares(Schuman et al., 2022).Typically, these application softwares include model construction tools, which are used to define the structure and parameters of neural networks, including neuron types, connection patterns. Simulators, which are used to simulate and debug neural network models on the chip, and optimization tools. For instance, the Neurogrid chip is paired with the Neurogrid Software Framework(Benjamin et al., 2014), allows users to specify neuronal models in the Python programming environment. The software framework for the BrainScaleS chip is the BrainScaleS-Software-Stack(Pehle et al., 2022). Its software interface supports training neural networks on the chip using the PyTorch framework. The TrueNorth chip typically utilizes a software framework called the TrueNorth Ecosystem which is developed by the TrueNorth native Corelet language(Akopyan et al., 2015). These application frameworks enable users to more conveniently construct, simulate, and optimize neural network models on the chips, facilitating efficient research and application development.
@@ 8.1.2 Image Generation/Denoising and Image/Video Reconstruction
Generation tasks are increasingly explored in SNNs. Comşa et al. (2021) introduced a directly trained spiking autoencoder to reconstruct images with high fidelity on MNIST and FMNIST datasets. Kamata et al. (2022) constructed a fully spiking variational autoencoder (FSVAE), generating images with competitive quality compared to conventional ANNs. Liu et al. (2023) proposed a Spiking-Diffusion model, outperforming the existing SNN-based generation models on several datasets. Castagnetti et al. (2023) developed an image denoising solution based on a directly trained spiking autoencoder, achieving a competitive signal-to-noise ratio on the Set12 dataset with significantly lower energy.
Visual information reconstruction is important for neuromorphic vision sensors, because humans cannot directly perceive visual scenes from events. Zhu and Tian (2023) provided a comprehensive review of visual reconstruction methods for events. Duwek et al. (2021) proposed a hybrid ANN-SNN model, accomplishing image reconstruction for simple scenes from N-MNIST and N-Caltech101 datasets. Zhu et al. (2021a) proposed an image reconstruction algorithm that combines DVS and Vidar signals, leveraging the high dynamic range of DVS to improve reconstruction effectiveness. Subsequently, they developed a deep SNN with an encoder-decoder structure for event-based video reconstruction, achieving performance comparable to ANN counterparts with only 0.05x energy consumption (Zhu et al., 2022).
@@ 8.1.3 Others
Besides the above-mentioned tasks, deep SNNs are also applied in some other computer vision tasks, including video action recognition (Panda and Srinivasa, 2018; Chakraborty and Mukhopadhyay, 2023; Wang et al., 2019; Zhang et al., 2022c), image segmentation (Liang et al., 2022; Parameshwara et al., 2021; Kim et al., 2022; Zhang et al., 2023a), optical flow estimation (Lee et al., 2020a; Kosta and Roy, 2023; Cuadrado et al., 2023), depth prediction (Rançon et al., 2022; Wu et al., 2022; Zhang et al., 2022b), point clouds processing (Zhou et al., 2020; Ren et al., 2023), human pose tracking (Zou et al., 2023), lip-reading (Bulzomi et al., 2023), emotion/expression recognition (Wang et al., 2022; Barchid et al., 2023), medical image classification (Shan et al., 2022; Qasim Gilani et al., 2023), and so on.

### v2

@@ Abstract
Spiking neural networks (SNNs) offer a promising energy-efficient alternative to artificial neural networks (ANNs), in virtue of their high biological plausibility, rich spatial-temporal dynamics, and event-driven computation. The direct training algorithms based on the surrogate gradient method provide sufficient flexibility to design novel SNN architectures and explore the spatial-temporal dynamics of SNNs. According to previous studies, the performance of models is highly dependent on their sizes. Recently, direct training deep SNNs have achieved great progress on both neuromorphic datasets and large-scale static datasets. Notably, transformer-based SNNs show comparable performance with their ANN counterparts. In this paper, we provide a new perspective to summarize the theories and methods for training deep SNNs with high performance in a systematic and comprehensive way, including theory fundamentals, spiking neuron models, advanced SNN models and residual architectures, software frameworks and neuromorphic hardware, applications, and future trends. The reviewed papers are collected in Awesome-Spiking-Neural-Networks.
@@ 3 Introduction
There are two mainstream pathways to obtain deep SNNs: ANN-to-SNN conversion and direct training through the surrogate gradient method. Firstly, in ANN-to-SNN conversion (Cao et al., 2015; Hunsberger and Eliasmith, 2015; Rueckauer et al., 2017; Bu et al., 2022; Meng et al., 2022; Wang et al., 2022), a pre-trained ANN is converted to an SNN by replacing the ReLU activation layers with spiking neurons and adding scaling operations like weight normalization and threshold balancing. This conversion process suffers from long converting time steps, which causes high computational consumption in practice. In addition, the converted SNNs obtained in this way are constrained by the original ANNs’ architecture and are hard to adapt to dynamic signal (DVS, DAVIS, ATIS data) processing. Thus, the direct exploration of the virtues of SNNs is limited in ANN-to-SNN conversion. Secondly, in the field of direct training, SNNs are unfolded over simulation time steps and trained with backpropagation through time (Lee et al., 2016; Shrestha and Orchard, 2018). Due to the non-differentiability of spiking neurons, the surrogate gradient method is employed for backpropagation (Lee et al., 2020b; Neftci et al., 2019; Fang et al., 2021a, b; Zhou et al., 2023c). On one hand, this direct training method can handle temporal data and also achieve decent performance on large-scale static datasets, with only a few time steps. On the other hand, it can provide sufficient flexibility for designing novel architectures specifically for SNNs and exploring the properties of SNNs directly. Therefore, the direct training method has received more attention recently.
@@ 4 Spiking Neuron Models
LIF (Leaky Integrate-and-Fire) neuron is one of the most commonly used neurons in SNNs (Zhou et al., 2023c, a, b), which is simple but retains biological characteristics (figure 1a). The dynamics of LIF are described as:
@@
Figure 1: (a) The scheme of a spiking neuron, of which the input and output are both binary spikes. (b) The sigmoid function approximates the Heaviside activation function of a spiking neuron, and its derivative can be utilized to calculate gradients during backpropagation.
@@ 5.2.1 Surrogate Gradient
Eq. 13 is the derivative of Eq. 12. In addition, SG could be the derivative of Sigmoid (figure 1b) (Zhou et al., 2023c, a, b), tanh (Guo et al., 2022a), or rectangular (Wu et al., 2018, 2019) functions, etc. To address the problem of gradient vanishing caused by surrogate gradient function with fixed parameters, Lian et al. proposed the Learnable Surrogate Gradient (LSG) (Lian et al., 2023), in which a learnable parameter is used to adjust the gradient-available interval.
@@ 5.2.2 Loss Function and Backpropagation
Figure 2: The backward of BPTT and OTTT.
Distinct from ANNs, there’s an additional dimension in SNNs, the temporal domain. For spiking neurons, the membrane potential in the current step depends on the membrane potential in the previous time step, that is, there is a time dependence. Thus, backpropagation in ANNs does not apply to SNNs. Backpropagation Through Time (BPTT) (Werbos, 1990; Bird and Polivoda, 2021), originally developed for recurrent neural networks (RNNs), is applied to SNNs due to their similar characteristics to those of RNNs. The combination of BPTT and surrogate gradient is the basic approach in SNNs. Spatio-temporal backpropagation (STBP) (Wu et al., 2018), proposed by Wu et al., takes the gradient update in both the spatial domain and temporal domain into account to train SNNs. However, the additional time dimension exposes BPTT and STBP to the problem of requiring a large amount of training memory and training time. Therefore, Xiao et al. proposed an online training through time (OTTT) (Xiao et al., 2022) algorithm derived from BPTT, which only requires constant training memory consumption agnostic to time steps, and reduces the significant memory costs compared to BPTT. The backward of BPTT and OTTT are shown in figure 2. Another efficient backpropagation method, Spatial Learning Through Time (SLTT) (Meng et al., 2023), ignores the unimportant routes in the computational graph during backpropagation, to reduce training memory consumption and training time. However, although OTTT and SLTT show better training memory consumption than BPTT, direct training high-performance SNNs are still dominated by the combination of BPTT and surrogate gradient, such as SGLFormer (Zhang et al., 2024), Spikformer (Zhou et al., 2023c), etc. Thus, it’s essential to investigate direct training methods offering both high effectiveness and efficiency.
@@ 6.1 Transformer-based Spiking Neural Networks
Biological realistic models tend to model neural networks with high biological plausibility to simulate the complex biological mechanism of the brain. It often lacks the consideration of computational efficiency and performance optimization on general application tasks. Traditional ANNs often prioritize task performance over biological realism and computational energy consumption. SNNs have great potential to own the characteristics of biological plausibility, low computational energy consumption, and high task performance simultaneously. Especially, several direct training Transformer-based SNNs have broken through 80% top-1 accuracy on ImageNet-1K, which instills great optimism in the application of SNNs.
6.2 Residual Architectures in Spiking Neural Networks
@@
6.3 Others
Besides the above-mentioned architectures, some other interesting research topics are also worthy of attention, such as Spiking RNN/LSTM, LSM, etc. Deep Liquid State Machine (LSM) (Wang and Li, 2016) explored the power of recurrent spiking networks and deep architectures. Soures and Kudithipudi (2019) proposed a novel deep LSM to capture dynamic information over multiple time-scales with a combination of randomly connected layers and unsupervised layers. Hamilton et al. (2019) demonstrated the nonlinear dynamics of spiking neurons can be used to implement low-level graph operations. Zhu et al. (2022b) proposed end-to-end Spiking Graph Convolutional Networks (GCNs) that integrate the embedding of GCNs with the biofidelity characteristics of SNNs. Bellec et al. (2020) and Bohnstingl et al. (2022) explored the architectures and online-training methods of recurrent spiking neural networks. Ren et al. (2023b) proposed a novel end-to-end point-based SNN architecture, which excels at processing sparse event cloud data, effectively extracting both global and local features through a singular-stage structure.
@@ 7.1 Software Frameworks for Training Spiking Neural Networks
One category of frameworks includes brain simulators such as NEURON (Hines and Carnevale, 1997) and Brian (Goodman and Brette, 2009), which not only enhance the scalability and computational efficiency of models but also encompass cognitive functions such as perception, decision-making, and reasoning. The SNNs constructed by these frameworks exhibit a high degree of biological plausibility, making them suitable for studying the functionalities of real neural systems. They support biologically interpretable learning rules such as Spike-Timing-Dependent Plasticity (STDP) (Bi and Poo, 1998), playing a significant role in advancing the field of neuroscience. However, these frameworks lack core computational functionalities required for deep learning, such as automatic differentiation, rendering them incapable of performing machine learning tasks. Another category of brain-inspired computing frameworks comprises deep spiking computation frameworks. Deep SNNs involve a substantial amount of matrix operations across spatial and temporal dimensions, a variety of neurons, neuromorphic datasets, and deployments on neuromorphic chips. The modeling and application processes are complex, and achieving high-performance acceleration is challenging. To address these issues, spiking deep learning frameworks need to support the construction, training, and deployment of deep SNNs, and be capable of acceleration based on spike operations. Frameworks such as BindsNET (Hazan et al., 2018), NengoDL (Rasmussen, 2019), SpykeTorch (Mozafari et al., 2019), Norse, SpyTorch, SNNTorch (Eshraghian et al., 2023) and SpikingJelly (Fang et al., 2023a) have been developed. They utilize simple spiking neurons to reduce computational complexity, making them suitable for machine learning research. Among them, BindsNet (Hazan et al., 2018) primarily focuses on machine learning and reinforcement learning; NengoDL (Rasmussen, 2019) converts ANNs to obtain deep SNNs but does not support direct training of SNNs using surrogate gradient methods; SpyTorch is a demonstrative framework that only provides basic surrogate gradient examples; SpyTorch (Mozafari et al., 2019) introduces a new type of surrogate gradient method named SuperSpike. These frameworks can implement some simple machine learning and reinforcement learning models, but they still lack deep learning capabilities for SNNs. Norse is attempting to introduce the sparse and event-driven characteristics of SNNs and supports many typical spiking neuron models. It is in the development stage and has not been officially released yet. SNNTorch supports some variants of online backpropagation algorithms that are more biologically plausible and support large-scale SNN computation. SpikingJelly (Fang et al., 2023a) is a full-stack toolkit for preprocessing neuromorphic datasets, building deep SNNs, optimizing their parameters, and deploying SNNs on neuromorphic chips, which shows remarkable extensibility and flexibility, enabling users to accelerate custom models at low costs through multilevel inheritance and semiautomatic code generation. In summary, the development of existing software frameworks is essentially in its early stages, and there is still a long way to go in terms of functionality enhancement and performance optimization.
@@ 7.2 Neuromorphic Hardware for Spiking Neural Networks
7.3 Software and Hardware Interplay
The deployment of algorithms for SNNs onto neuromorphic chips typically requires certain software frameworks. The computational software frameworks mentioned in Section 7.1 usually support simulations on mainstream CPUs and GPUs, without clear mention of deployment on neuromorphic chips. Meanwhile, among the previously mentioned neuromorphic chips in Section 7.2, only about 27 percent of them are connected to application software packages (Schuman et al., 2022). Typically, these application software packages contain model construction tools, simulators, and optimization tools. Model construction tools are used to define the structure and parameters of neural networks, including neuron types, connection patterns. Simulators are applied to simulate and debug neural network models on the chip. Optimization tools are adopted to train the network parameters and optimize its performance. Here are some typical examples. The Neurogrid chip is paired with the Neurogrid Software Framework (Benjamin et al., 2014), allowing users to specify neuronal models in the Python programming environment. The software framework for the BrainScaleS chip is the BrainScaleS-Software-Stack (Pehle et al., 2022), which supports training neural networks on the chip using the PyTorch framework. The IBM TrueNorth chip typically utilizes a software framework called the TrueNorth Ecosystem, which is developed by the TrueNorth native Corelet language (Akopyan et al., 2015). IBM NorthPole (Modha et al., 2023) is a brain-inspired memory-near-compute chip with a software development kit, but this chip does not emulate spiking communication. Tianjic’s software toolchain supports both ANN-to-SNN conversion and direct training for SNNs, and supports automatically transforming a pretrained model into an equivalent network that meets the Tianjic hardware constraints for non-spiking ANNs (Pei et al., 2019). The latest Darwin3 builds a specialized instruction set architecture (ISA) (Ma et al., 2024), which is close to machine code tailored for efficient neuromorphic computing. These software frameworks enable users to conveniently construct, simulate, and optimize neural network models on neuromorphic chips, facilitating efficient research and application development.
@@ 8.1.2 Image Generation/Denoising and Image/Video Reconstruction
Generation tasks are increasingly explored in SNNs. Comşa et al. (2021) introduced a directly trained spiking autoencoder to reconstruct images with high fidelity on MNIST and FMNIST. Kamata et al. (2022) constructed a fully spiking variational autoencoder (FSVAE), generating images with competitive quality compared to conventional ANNs. Liu et al. (2023) proposed a Spiking-Diffusion model, outperforming the existing SNN-based generation models on several datasets. Castagnetti et al. (2023) developed an image denoising solution based on a directly trained spiking autoencoder, achieving a competitive signal-to-noise ratio on the Set12 dataset with significantly lower energy.
Visual information reconstruction is important for neuromorphic vision sensors, because humans cannot directly perceive visual scenes from events. Zhu and Tian (2023) provided a comprehensive review of visual reconstruction methods for events. Duwek et al. (2021) proposed a hybrid ANN-SNN model, accomplishing image reconstruction for simple scenes from N-MNIST and N-Caltech101 datasets. Zhu et al. (2021a) proposed an image reconstruction algorithm that combines DVS and Vidar signals, leveraging the high dynamic range of DVS to improve reconstruction effectiveness. Subsequently, they developed a deep SNN with an encoder-decoder structure for event-based video reconstruction, achieving performance comparable to ANN counterparts with only 0.05x energy consumption (Zhu et al., 2022a).
@@ 8.1.3 Others
Besides the above-mentioned tasks, deep SNNs are also applied in some other computer vision tasks, including video action recognition (Panda and Srinivasa, 2018; Chakraborty and Mukhopadhyay, 2023; Wang et al., 2019; Zhang et al., 2022c), image segmentation (Liang et al., 2022; Parameshwara et al., 2021; Kim et al., 2022; Zhang et al., 2023a), optical flow estimation (Lee et al., 2020a; Kosta and Roy, 2023; Cuadrado et al., 2023), depth prediction (Rançon et al., 2022; Wu et al., 2022; Zhang et al., 2022b), point clouds processing (Zhou et al., 2020; Ren et al., 2023a), human pose tracking (Zou et al., 2023), lip-reading (Bulzomi et al., 2023), emotion/expression recognition (Wang et al., 2022; Barchid et al., 2023), medical image classification (Shan et al., 2022; Qasim Gilani et al., 2023), and so on.